# Spike No More: Stabilizing the Pre-training of Large Language Models

### v2

@@ Abstract
Loss spikes often occur during pre-training of large language models. The spikes degrade the performance of large language models and sometimes ruin the pre-training. Since the pre-training needs a vast computational budget, we should avoid such spikes. To investigate the cause of loss spikes, we focus on gradients of internal layers. Through theoretical analyses, we reveal two causes of the exploding gradients, and provide requirements to prevent the explosion. In addition, we propose a method to satisfy the requirements by combining the initialization method and a simple modification to embeddings. We conduct various experiments to verify our theoretical analyses empirically. Experimental results indicate that the combination is effective in preventing spikes during pre-training.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLMs) have been fundamental assets for various applications (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023). Increasing the number of parameters in (neural) language models and the number of training data usually leads to better LLMs (Kaplan et al., 2020). Consequently, the pre-training requires a vast budget, and thus, minimizing the risk of failure of the pre-training is a paramount concern.
Despite their widespread use as the foundational architecture for LLMs, a comprehensive theoretical understanding of Transformers (Vaswani et al., 2017) has not yet been achieved. One of the crucial unresolved questions is the reason for the frequent occurrence of pre-training failures in Transformer-based LLMs due to spikes in loss values (loss spike) that can lead to catastrophic divergence (Chowdhery et al., 2022) as illustrated in Vanilla in Figure 1. While several empirical strategies have been proposed to mitigate this issue (Chowdhery et al., 2022; Le Scao et al., 2022; Zeng et al., 2023), the absence of theoretical justification for these methods casts unclear on their generalizability to other situations, such as varying sizes of model parameters.
@@
In this research, we provide theoretical analyses focusing on the loss spike issue during LLM pre-training. We identify two primary factors that lead to the issue: (1) the rapid amplification of the norm of the shortcut part in the residual connection during forward propagation and (2) the intensification of gradients before and after layer normalizations (LNs) (Ba et al., 2016) in each layer. Our research also demonstrates that the impact of each factor is closely related to the chosen parameter initialization technique. We provide evidence that initializing parameters using a typical methodology (He et al., 2015) results in gradient explosion attributable to the shortcut. Conversely, a popular initialization method in LLM pre-training, which reduces certain parameters to smaller values, promotes the exploding gradients due to LNs although it mitigates the gradient explosion arising from the shortcut. To prevent both exploding gradients, we describe a requirement and introduce a simple modification to satisfy the requirement.
To substantiate our theoretical analyses, we conduct a series of empirical experiments. The results show that a method satisfying the requirement hinders the occurrence of loss spikes and explosions. This contrasts with methods that fail to satisfy the requirement, which remain susceptible to the exploding gradients even when previously advocated as an empirical solution. Additionally, we demonstrate that a method satisfying the requirement enables LLMs to be pre-trained with a comparatively larger learning rate, culminating in superior performance outcomes.
@@ 2.1 Pre-LN Transformer
This paper mainly focuses on the neural architecture used in the GPT series (Radford et al., 2018, 2019; Brown et al., 2020). They use the Pre-LN Transformer (Xiong et al., 2020) which is the de facto standard architecture in recent implementations of Transformers because the architecture is more stable than the original Transformer architecture when we stack many layers (Xiong et al., 2020; Liu et al., 2020; Takase et al., 2023). Let $x\in\mathbb{R}^{d}$ be an input of a layer of the Transformer, where $d$ denotes the dimension of the layer. The layer outputs $y$ with the following equations:
@@
$x^{\prime}$ $=x+\mathrm{Attention}(\mathrm{LN}(x)),$ (2)
@@
where $\mathrm{LN}$ is the layer normalization function. We call the first terms in Equations (1) and (2), i.e., $x$ and $x^{\prime}$, shortcut. In addition, $\mathrm{FFN}$ and $\mathrm{Attention}$ are defined as follows [^1] To simplify equations, we omit bias terms.:
$\mathrm{FFN}(x)$ $=\mathcal{F}(xW_{1})W_{2},$ (3)
$\mathrm{Attention}(x)$ $=\mathrm{concat}(\mathrm{head}_{1}(x),...,\mathrm{head}_{h}(x))W_{O},$ (4)
$\mathrm{head}_{i}(x)$ $=\mathrm{softmax}\left(\frac{xW_{Qi}(XW_{Ki})^{\mathrm{T}}}{\sqrt{d_{head}}}\right)XW_{Vi},$ (5)
@@
where $\mathcal{F}$ is an activation function, $\mathrm{concat}$ concatenates input vectors, $\mathrm{softmax}$ applies the softmax function to an input vector, and $W_{1}\in\mathbb{R}^{d\times d_{ffn}}$, $W_{2}\in\mathbb{R}^{d_{ffn}\times d}$, $W_{Qi}\in\mathbb{R}^{d\times d_{head}}$, $W_{Ki}\in\mathbb{R}^{d\times d_{head}}$, $W_{Vi}\in\mathbb{R}^{d\times d_{head}}$, and $W_{O}\in\mathbb{R}^{d\times d}$ are parameter matrices, and $d_{ffn}$ and $d_{head}$ are the internal dimensions of FFN and multi-head attention sub-layers, respectively. In addition, we pack the sequence of input vectors into a matrix as $X$ to compute the self-attention.
We discuss the difference from the original Transformer, i.e., Post-LN Transformer in Appendix D and using Root Mean Square layer normalization (RMSNorm) instead of LN in Section 5.3.
2.2 Initialization

### v3

@@ Abstract
Loss spikes often occur during pre-training of large language models. The spikes degrade the performance of large language models and sometimes ruin the pre-training. Since the pre-training needs a vast computational budget, we should avoid such spikes. Based on the assumption that the loss spike is caused by the sudden growth of the gradient norm, we explore factors to keep the gradient norm small through an analysis of the spectral norms of the Jacobian matrices for the sub-layers. Our findings suggest that stabilizing the pre-training process requires two conditions: small sub-layers and large shortcut. We conduct various experiments to empirically verify our theoretical analyses. Experimental results demonstrate that methods satisfying the conditions effectively prevent loss spikes during pre-training.
@@ 1 Introduction
Large language models (LLMs) have been fundamental assets for various applications (Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023). Increasing the number of parameters in (neural) language models and the number of training data usually leads to better LLMs (Kaplan et al., 2020). Consequently, pre-training requires a vast budget, and thus, minimizing the risk of failure of the pre-training is a paramount concern.
Despite their widespread use as the foundational architecture for LLMs, a comprehensive theoretical understanding of Transformers (Vaswani et al., 2017) has not yet been achieved. One of the crucial unresolved questions is the reason for the frequent occurrence of pre-training failures in Transformer-based LLMs due to spikes in loss values (loss spike) that can lead to catastrophic divergence (Chowdhery et al., 2022) as illustrated in Vanilla in Figure 1. While several empirical strategies have been proposed to mitigate this problem (Chowdhery et al., 2022; Le Scao et al., 2022; Zeng et al., 2023), the absence of theoretical justification for these methods casts unclear on their generalizability to other situations, such as varying sizes of model parameters.
In this research, we provide theoretical analyses focusing on the loss spike problem during LLM pre-training. We identify the upper bound of the gradient norms for the Transformer-based LLMs through analyses on the spectral norms of the Jacobian matrices for the sub-layers. If the upper bound is large, the gradients may spike suddenly, and we assume that this phenomenon causes the loss spike. Then, we indicate that the upper bound is large in the typical setting, such as the widely used implementation, Megatron-LM (Shoeybi et al., 2020), and thus, the loss spike is likely to occur. In addition, to make the upper bound sufficiently small, we introduce two conditions: (1) initializing the parameters of sub-layers with a small value and (2) making the standard deviation of each embedding close to $1$. The former condition can be satisfied by the widely used initialization method for LLMs (Shoeybi et al., 2020; Le Scao et al., 2022; Biderman et al., 2023). On the other hand, the latter condition was satisfied in the original Transformer by scaling embeddings (Vaswani et al., 2017), but such scaling is missing from recent implementations. To sum up, through theoretical analyses, we re-evaluate several previous techniques in terms of the stabilization of LLM pre-training.
Building on our theoretical analysis, we further substantiate our claims through a series of empirical experiments, which provides a clear distinction between effective and ineffective methods over different training scenarios. Our results demonstrate that methods satisfying the conditions avoid the occurrence of loss and gradient spikes. In contrast, methods that fail to meet these conditions remain susceptible to gradient spikes, even when previously recommended as empirical solutions of the loss spike problem. Furthermore, we demonstrate that a method satisfying the conditions enables LLMs to be pre-trained with a comparatively larger learning rate, leading to superior performance outcomes.
@@ 2.1 Pre-LN Transformer
This paper mainly focuses on the neural architecture used in the GPT series (Radford et al., 2018; 2019; Brown et al., 2020). They use the Pre-LN Transformer (Xiong et al., 2020), which is the de facto standard architecture in recent implementations of Transformers because the training with the architecture is more stable than the original Transformer architecture when we stack many layers (Xiong et al., 2020; Liu et al., 2020; Takase et al., 2023). Let $x\in\mathbb{R}^{d}$ be an input of a layer of the Transformer, where $d$ denotes the dimension of the layer. The layer outputs $y$ with the following equations:
@@
$x^{\prime}$ $=x+\mathrm{Attn}(\mathrm{LN}(x)),$ (2)
@@
where $\mathrm{LN}$ is the layer normalization function [^1] We discuss the difference from the architecture using Root Mean Square layer normalization (RMSNorm) (Zhang & Sennrich, 2019) instead of LN in Appendix D, and the original Transformer architecture, i.e., Post-LN Transformer in Appendix J.. We call the first terms in Equations (1) and (2), i.e., $x$ and $x^{\prime}$, shortcut. In addition, the feed-forward network ($\mathrm{FFN}$) and multi-head self-attention ($\mathrm{Attn}$) are defined as follows [^2] To simplify equations, we omit bias terms.:
$\mathrm{FFN}(x)$ $=W_{2}(\mathcal{F}(W_{1}\ x)),$ (3)
$\mathrm{Attn}(x)$ $=W_{O}(\mathrm{concat}(\mathrm{head}_{1}(x),...,\mathrm{head}_{h}(x))),$ (4)
$\mathrm{head}_{i}(x)$ $=\mathrm{softmax}\left(\frac{(W_{Qi}\ x)^{\mathrm{T}}(W_{Ki}\ X)}{\sqrt{d_{\mathrm{head}}}}\right)(W_{Vi}\ X)^{\mathrm{T}},$ (5)
@@
where $\mathcal{F}$ is an activation function, $\mathrm{concat}$ concatenates input vectors, $\mathrm{softmax}$ applies the softmax function to an input vector, and $W_{1}\in\mathbb{R}^{d_{\mathrm{ffn}}\times d}$, $W_{2}\in\mathbb{R}^{d\times d_{\mathrm{ffn}}}$, $W_{Qi}\in\mathbb{R}^{d_{\mathrm{head}}\times d}$, $W_{Ki}\in\mathbb{R}^{d_{\mathrm{head}}\times d}$, $W_{Vi}\in\mathbb{R}^{d_{\mathrm{head}}\times d}$, and $W_{O}\in\mathbb{R}^{d\times d}$ are parameter matrices, and $d_{\mathrm{ffn}}$ and $d_{\mathrm{head}}$ are the internal dimensions of FFN and multi-head self-attention sub-layers, respectively. In addition, we pack the sequence of input vectors into a matrix as $X\in\mathbb{R}^{d\times L}$, where $L$ is the input sequence length, to compute the self-attention.
2.2 Gradients of Pre-LN Transformers
Let $\mathcal{L}$ be the loss function of the $N$ layered Pre-LN Transformer and $J_{n}$ be the Jacobian matrix of the $n$-th layer. We can calculate the gradient of $\mathcal{L}$ using the relations in Equations (1) and (2) as:
$\frac{\partial\mathcal{L}}{\partial x_{1}}=\frac{\partial\mathcal{L}}{\partial y_{N}}\prod_{n=1}^{N-1}J_{n}=\frac{\partial\mathcal{L}}{\partial y_{N}}\prod_{n=1}^{N-1}\left(\frac{\partial y_{n}}{\partial x^{\prime}_{n}}\frac{\partial x^{\prime}_{n}}{\partial x_{n}}\right),\quad\mbox{where}\quad J_{n}=\frac{\partial y_{n}}{\partial x_{n}}$ $=\frac{\partial y_{n}}{\partial x^{\prime}_{n}}\frac{\partial x^{\prime}_{n}}{\partial x_{n}}.$ (6)
Using the submultiplicativity of the spectral norm, i.e., $\|AB\|_{2}\leq\|A\|_{2}\|B\|_{2}$, and Equation (6), we can derive an upper bound of the norm of the gradient of $\mathcal{L}$ as:
$\bigg{\|}\frac{\partial\mathcal{L}}{\partial x_{1}}\bigg{\|}_{2}=\bigg{\|}\frac{\partial\mathcal{L}}{\partial y_{N}}\prod_{n=1}^{N-1}\frac{\partial y_{n}}{\partial x^{\prime}_{n}}\frac{\partial x^{\prime}_{n}}{\partial x_{n}}\bigg{\|}_{2}$ $\leq\bigg{\|}\frac{\partial\mathcal{L}}{\partial y_{N}}\bigg{\|}_{2}\prod_{n=1}^{N-1}\bigg{\|}\frac{\partial y_{n}}{\partial x^{\prime}_{n}}\bigg{\|}_{2}\bigg{\|}\frac{\partial x^{\prime}_{n}}{\partial x_{n}}\bigg{\|}_{2}.$ (7)
Thus, we can estimate the upper bound of the gradient norm of $\mathcal{L}$ by analyzing the spectral norms of the Jacobian matrices for the FFN layer and the self-attention layer, namely, $\|\frac{\partial y_{n}}{\partial x^{\prime}_{n}}\|_{2}$ and $\|\frac{\partial x^{\prime}_{n}}{\partial x_{n}}\|_{2}$.
2.3 Motivation to Suppress the Upper Bound
In our preliminary experiments, when the gradient norms grow suddenly during LLM pre-training, we observe that the loss spike problem is likely to occur. Thus, we assume that we can prevent the loss spike problem by maintaining the gradient norm small. To prevent the growth of the gradient norm, we explore the way to suppress the upper bound described by Equation (7). To suppress the upper bound, we analyze the Jacobian matrices to find a factor to control the upper bound in the following sections, and then, provide two conditions: small sub-layers and large shortcut. We verify our assumption and theoretical analyses through experiments on LLM pre-training.
3 Analyses on Gradients of Sub-Layers
For the theoretical analyses in this section, we employ the following assumption:
Assumption 1.
Let $x$ and $x^{\prime}$ be the input and intermediate vectors of each layer. Moreover, let $W_{*}$ denote the model parameter matrix in each layer. We assume that $x$, $x^{\prime}$, and $W$ for all layers follow a normal distribution with a mean of 0, i.e., $\mu=0$.
This assumption is valid when we initialize parameters with the normal distribution, the number of heads in Equation (4) is $1$, and $\mathcal{F}$ is an identity function. Empirically, the outputs of each sub-layer are close to the normal distribution as illustrated in Appendix F.
3.1 Jacobian Matrix of FFN
Based on Equation (1), $\|\frac{\partial y_{n}}{\partial x^{\prime}_{n}}\|_{2}$ in Equation (7) can be rewritten as:
$\bigg{\|}\frac{\partial y}{\partial x^{\prime}}\bigg{\|}_{2}$ $=\bigg{\|}\frac{\partial(x^{\prime}+\mathrm{FFN}(\mathrm{LN}(x^{\prime})))}{\partial x^{\prime}}\bigg{\|}_{2}=\bigg{\|}I+\frac{\partial(\mathrm{FFN}(\mathrm{LN}(x^{\prime})))}{\partial x^{\prime}}\bigg{\|}_{2}.$ (8)
We can then derive an upper bound of $\|\frac{\partial y_{n}}{\partial x^{\prime}_{n}}\|_{2}$ by applying the subadditivity, i.e., $\|A+B\|_{2}\leq\|A\|_{2}+\|B\|_{2}$, and submultiplicativity properties of the spectral norm as follows:
$\bigg{\|}\frac{\partial y}{\partial x^{\prime}}\bigg{\|}_{2}$ $\leq 1+\bigg{\|}\frac{\partial\mathrm{FFN}(\mathrm{LN}(x^{\prime}))}{\partial\mathrm{LN}(x^{\prime})}\bigg{\|}_{2}\bigg{\|}\frac{\partial\mathrm{LN}(x^{\prime})}{\partial x^{\prime}}\bigg{\|}_{2}.$ (9)
The right-hand side of this inequality indicates that we can estimate the upper bound of $\|\frac{\partial y}{\partial x^{\prime}}\|_{2}$ by separately computing the spectral norms of Jacobian matrices for $\mathrm{FFN}$ and $\mathrm{LN}$.
Regarding the $\mathrm{FFN}$ part, we assume that the activation function $\mathcal{F}$ is an identity function [^3] Appendix G discusses the case where we use the $\mathrm{ReLU}$ as the activation function, which results in the same conclusion with a slightly different coefficient. to simplify the discussion. Under this assumption, the following equation holds:
$\bigg{\|}\frac{\partial\mathrm{FFN}(\mathrm{LN}(x^{\prime}))}{\partial\mathrm{LN}(x^{\prime})}\bigg{\|}_{2}=\|W_{1}W_{2}\|_{2}.$ (10)
Therefore, we can straightforwardly derive the relation $\|W_{1}W_{2}\|_{2}\leq\|W_{1}\|_{2}\|W_{2}\|_{2}$ from the submultiplicativity of the spectral norm. Furthermore, let $\sigma_{1}$ and $\sigma_{2}$ be the standard deviations of $W_{1}$ and $W_{2}$, respectively. From Assumption 1, the spectral norms of $W_{1}$ and $W_{2}$ are obtained by their standard deviations and dimensions (Vershynin, 2018), i.e., $\|W_{1}\|_{2}\approx\sigma_{1}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})$ and $\|W_{2}\|_{2}\approx\sigma_{2}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})$. Finally, we can express an upper bound of the spectral norms of the Jacobian matrices for $\mathrm{FFN}$ as the following inequality:
$\bigg{\|}\frac{\partial\mathrm{FFN}(\mathrm{LN}(x^{\prime}))}{\partial\mathrm{LN}(x^{\prime})}\bigg{\|}_{2}\leq\sigma_{1}\sigma_{2}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})^{2},$ (11)
where the right-hand side has the relation $\sigma_{1}\sigma_{2}(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})^{2}=\|W_{1}\|_{2}\|W_{2}\|_{2}$.
Next, regarding the $\mathrm{LN}$ part, the Jacobian matrix of LN can be written as:
$\frac{\partial\mathrm{LN}(x^{\prime})}{\partial x^{\prime}}=\frac{\sqrt{d}}{\|x^{\prime}\|_{2}}\bigg{(}I-\frac{x^{\prime}x^{\prime\top}}{\|x^{\prime}\|_{2}^{2}}\bigg{)}$ $=\frac{\sqrt{d}}{\sigma_{x^{\prime}}\sqrt{d}}\bigg{(}I-\frac{x^{\prime}x^{\prime\top}}{\sigma_{x^{\prime}}^{2}d}\bigg{)}=\frac{1}{\sigma_{x^{\prime}}}\bigg{(}I-\frac{zz^{\top}}{d}\bigg{)}.$ (12)
The leftmost equation appears in the proof by Xiong et al. (2020). The second equation uses $\|x^{\prime}\|_{2}=\sigma_{x^{\prime}}\sqrt{d}$, which can be obtained based on Assumption 1. The last equation is derived from the well-known formula of $z=(x^{\prime}-\mu_{x^{\prime}})/\sigma_{x^{\prime}}$, which converts a normal distribution, $x^{\prime}$, to the standard normal distribution $z$, where $\mu_{x^{\prime}}=0$ in Assumption 1.
We consider the variance ($\mathrm{var}$) of each element in the matrix $zz^{\top}$. Since $z_{i}z_{i}$ follows $\mathcal{X}^{2}$ with $1$ degree of freedom, and $z_{i}z_{j}(i\neq j)$ is the multiplication of two independent values following the standard normal distribution, the variances are as follows:
$\mathrm{var}(z_{i}z_{j})=\begin{cases}1\ \ \text{if $i\neq j$}\\ 2\ \ \text{otherwise}\end{cases}.$ (13)
Equation (13) indicates that $\frac{zz^{\top}}{d}\approx 0$ in LLMs due to $d\gg 1$. Therefore, the spectral norm of the Jacobian matrix of LN can be written as:
$\bigg{\|}\frac{\partial\mathrm{LN}(x^{\prime})}{\partial x^{\prime}}\bigg{\|}=\frac{1}{\sigma_{x^{\prime}}},\quad\mbox{where}\quad\frac{\partial\mathrm{LN}(x^{\prime})}{\partial x^{\prime}}=\frac{1}{\sigma_{x^{\prime}}}I.$ (14)
Finally, Equation (9) can be rewritten by substituting Equations (11) and (14) as:
$\bigg{\|}\frac{\partial y}{\partial x^{\prime}}\bigg{\|}_{2}$ $\leq 1+\frac{\sigma_{1}\sigma_{2}}{\sigma_{x^{\prime}}}C_{\mathrm{ffn}},$ (15)
where $C_{\mathrm{ffn}}=(\sqrt{d}+\sqrt{d_{\mathrm{ffn}}})^{2}$ for the simplification.
According to the discussion in Section 2.3 and Equation (15), the standard deviations, $\sigma_{1}$ and $\sigma_{2}$, of $W_{1}$ and $W_{2}$, respectively, should be sufficiently small, and the standard deviation, $\sigma_{x^{\prime}}$, of the shortcut, $x^{\prime}$, should satisfy $\sigma_{1}\sigma_{2}\ll\sigma_{x^{\prime}}$ in order to keep the upper bound small.
3.2 Jacobian Matrix of Self-Attention
Similar to $\mathrm{FFN}$, we can rewrite $\|\frac{\partial x^{\prime}}{\partial x}\|_{2}$ in Equation (7) by using Equation (2) as:
$\bigg{\|}\frac{\partial x^{\prime}}{\partial x}\bigg{\|}_{2}$ $=\bigg{\|}\frac{\partial(x+\mathrm{Attn}(\mathrm{LN}(x)))}{\partial x}\bigg{\|}_{2}=\bigg{\|}I+\frac{\partial(\mathrm{Attn}(\mathrm{LN}(x)))}{\partial x}\bigg{\|}_{2}.$ (16)
We can then derive an upper bound of $\|\frac{\partial x^{\prime}}{\partial x}\|_{2}$ by applying the subadditivity and submultiplicativity of the spectral norm, namely:
$\bigg{\|}\frac{\partial x^{\prime}}{\partial x}\bigg{\|}_{2}$ $\leq 1+\bigg{\|}\frac{\partial\mathrm{Attn}(\mathrm{LN}(x))}{\partial\mathrm{LN}(x)}\bigg{\|}_{2}\bigg{\|}\frac{\partial\mathrm{LN}(x)}{\partial x}\bigg{\|}_{2}.$ (17)
Therefore, to estimate the upper bound of $\|\frac{\partial x^{\prime}}{\partial x}\|_{2}$, we compute the spectral norms of the Jacobian matrices for $\mathrm{Attn}$ and $\mathrm{LN}$.
Let $Z(\cdot)=\mathrm{concat}(\mathrm{head}_{1}(\cdot),...,\mathrm{head}_{h}(\cdot)))$ and let $J^{Z}$ be the Jacobian of the $Z(\cdot)$ [^4] We discuss the detail of $J^{Z}$ in Appendix I., we can rewrite the spectral norm of the Jacobian matrix of Attn as:
$\bigg{\|}\frac{\partial\mathrm{Attn}(\mathrm{LN}(x))}{\partial\mathrm{LN}(x)}\bigg{\|}_{2}=\bigg{\|}\frac{\partial W_{O}Z(\mathrm{LN}(x))}{\partial Z(\mathrm{LN}(x))}\frac{\partial Z(\mathrm{LN}(x))}{\partial\mathrm{LN}(x)}\bigg{\|}_{2}=\|W_{O}J^{Z}\|_{2}.$ (18)
Therefore, we can straightforwardly derive the relation $\|W_{O}J^{Z}\|_{2}\leq\|W_{O}\|_{2}\|J^{Z}\|_{2}$ from the submultiplicativity of the spectral norm.
Let $\sigma_{O}$ be the standard deviation of $W_{O}$. The relation $\|W_{O}\|_{2}\approx\sigma_{O}(2\sqrt{d})$ is derived from Assumption 1. We assign this value to Equation (18) and obtain the following inequality:
$\bigg{\|}\frac{\partial\mathrm{Attn}(\mathrm{LN}(x))}{\partial\mathrm{LN}(x)}\bigg{\|}_{2}\leq\sigma_{O}(2\sqrt{d})\|J^{Z}\|_{2}.$ (19)
Therefore, we can rewrite Equation (17) by substituting Equations (14) and (19) as follows:
$\bigg{\|}\frac{\partial x^{\prime}}{\partial x}\bigg{\|}_{2}$ $\leq 1+\frac{\sigma_{O}}{\sigma_{x}}C_{\mathrm{Attn}},$ (20)
where $C_{\mathrm{Attn}}=(2\sqrt{d})\|J^{Z}\|_{2}$ for the simplification.
Thus, similar to the discussion at the end of Section 3.1, the standard deviation, $\sigma_{O}$, of $W_{O}$ should be small and the standard deviation, $\sigma_{x}$, of the shortcut, $x$, should satisfy $\sigma_{O}\ll\sigma_{x}$ in order to keep the upper bound small.
4 Conditions to Avoid Spikes